A company has an income of $100,000 before paying taxes and a bonus. The bonus B is to be 20% of the income after deducting income taxes T but before deduction the bonus. So
B=0.20(100,000-T)
Because the bonus is a deductible expense, the amount of income tax T at a 40% rate is 40% of the income after deducting the bonus. So
T=0.40(100,000-B)
*Solve the system algebraically to find the bonus and the amount of tax.

Answer by stanbon(75887)

T=0.40(100,000-B)
*Solve the system algebraically to find the bonus and the amount of tax.
T=0.40(100,000-B)
*Solve the system algebraically to find the bonus and the amount of tax.
--------------
Tax = 40,000 - 0.4B
Bonus = (40,000-Tax)/0.40
============================
Cheers,
Stan H.